The box stands on a flat part of the truck moving at a speed of 15 m / s on the road. The driver abruptly applies the brakes and the vehicle stops after 38 m. Assuming that the truck acceleration is constant, determine the minimum value of the coefficient of friction that prevents the box from sliding on the ground.


1
Expert's answer
2021-03-24T19:43:41-0400
"v^2=2as\\v^2=2gs\mu\\15^2=2(9.8)(38)\mu\\ \mu=0.30"